# M-SET: Multi-Drone Swarm Intelligence Experimentation with Collision Avoidance Realism

### v1

@@ III-A4 Collision detection and collision avoidance
It detects all possible in-flight collisions, i.e., intersections of flight paths in multi-drone missions, and then minimizes the likelihood of collisions during the path planning. This requires an intelligent path planning algorithm, making distributed sensing cost-effective and safe within various scenarios. The algorithm detects intersections of flight paths, and augments the navigation and sensing plans selected via swarm intelligence to prevent the potential collisions, minimizing in-flight risks, i.e., the traveling distance at high risk of collisions. Fig 1 illustrates three typical types of collisions considered in this paper (cross, parallel, and destination-occupied).
@@ III-B Architecture overview
Fig. 2 illustrates an overview of M-SET architecture. The core of M-SET lies in two software approaches. One is the decentralized multi-agent collective learning method, named Economic Planning and Optimized Selections (EPOS) [6, 7, 22]. It generates for each agent a finite number of discrete navigation and sensing options, each with an estimated power consumption: the possible plans and their cost respectively. Plan generation is performed using the drones specifications (weight, propeller and battery parameters), and grid cell information about the sensing environment. Then the agents interact iteratively in a bottom-up and top-down fashion over a tree communication structure to make a selection such that all choices together add up to maximize the sensing quality (matching). However, these selected plans have no information about the potential collisions, and thus the path planning algorithm for collision avoidance is required to augment selected plans. This testbed uses the artificial potential field algorithm [5] to generate attractive forces towards the destinations and repulsive forces between drones. These forces guide the drones to their target cells for sensing according to the selected plan, while also avoiding potential collisions. Finally, the algorithm produces the output plans containing the X/Y coordinates and sensing (hovering) duration at each cell. They are then executed by drones in hardware, which records the energy consumption, the collected sensor data and collision risk distance during the sensing mission. Therein, collision risk distance implies the specific distance at which the risk of collisions becomes significant, illustrated by the path length within the shaded area shown in Fig. 1.
@@ IV-A Hardware UAVs for sensing
M-SET uses the Crazyflie 2.1 because of its size, weight and accessibility [^3] https://www.bitcraze.io/products/crazyflie-2-1/. The drone has 27g weight, 47mm propeller length, a battery capacity of 250mAh (LiPo battery) and it is configured to fly with an average ground speed of 0.1m/s. Based on this information, the power consumption can be estimated. Crazyflie can be programmed in Python with the support of API in Bitcraze [^4] https://github.com/bitcraze/crazyflie-lib-python. This is further expanded by the Crazyswarm API [^5] https://github.com/USC-ACTLab/crazyswarm, which provides additional functionality to control a swarm of Crazyflies. The drone can be mounted by multiple hardware decks to support different functions, such as positioning (with the support of lighthouse base stations), camera recording, and wireless charging. Due to maximum weight limitation of Crazyflie, two types of drones are assembled, one is for testing wireless charging, see Fig. 3(a), and the other is for collecting sensor data, see Fig. 3(b). The data-collection drone with an ultra low power 320×320 grayscale camera is equipped with a tiny mirror to take videos of the ground.
@@ IV-B Setting up an indoor sensing environment
A 75 inch screen is set on the ground as an indoor environment to emulate the outdoor sensing environments. As shown in Fig. 4(a) and Fig. 4(b), the screen is divided into 2x3 square grid cells, each with an area of 55x47cm. Each cell can be sensed effectively from an altitude of 50cm (i.e., the vehicles can be clearly observed). Other accessory equipment is set, including lighthouse base stations and wireless chargers. To show the significant and broad impact of M-SET on a transportation scenario, the screen displays the video of the traffic flow of vehicles recorded from satellites or other simulators (e.g., simulations of urban mobility  [^6] https://eclipse.dev/sumo/). In this scenario, a higher sensing requirement in a cell represents a high urgency for traffic monitoring over this area, which requires drones to hover a longer time to measure accurately. M-SET uses a real-world dataset of vehicle trajectories named pNEUMA [^7] https://open-traffic.epfl.ch/ collected by a swarm of drones in the congested downtown area of Athens, Greece [8].
@@ IV-C Swarm intelligence using collective learning
Figure 5: An example of collision avoidance using artificial potential field algorithm: (a) The blue drone is the target drone attracted by the destination, whereas the red drone is the obstacle drone repelling the target drone. The vectors influenced by both attractive and repulsive forces point towards the navigation of the target drone. (b) Since drone 2 has higher priority and stronger repulsive force than drone 1, drone 1 is pushed out of the target cell and "wait" until drone 2 passes by, and makes a "turn" when traveling to the next cell.
@@ IV-D Collision avoidance using artificial potential field
M-SET applies an artificial potential field algorithm [5] to path planning of drones, which creates force to repel drones from obstacles and attract them towards their destinations. As shown in Fig. 5, a Potential Fields Grid (PFG) in this implementation is created for each drone (named as the target drone) per timestamp. It is a 2D-grid of vectors, where each vector points in the direction that the target drone should fly at that position per timestamp [5]. There are two types of PFG: attractive PFG and repulsive PFG, which coexist to balance the drone navigation. The attractive PFG generates forces that pull the drone towards its target destination. The repulsive PFG creates forces that push the drone away from obstacles (e.g., the other drones). These repulsive forces are stronger when the target drone is closer to an obstacle. After the summation of all types of forces, the drone navigates towards its destination while avoiding obstacles. Thus, given $I$ vectors, a vector with index of $i$ at timestamp $t$, $i\geq 1$, $t\geq 1$, consists of two components: attractive component $V^{\mathsf{a}}$ and repulsive component $V^{\mathsf{r}}$, formulated as follows:
$V_{i}(t)=V^{\mathsf{a}}_{i}(t)+\sum_{j\in\mathcal{N}^{-}}V^{\mathsf{r}}_{i,j}(t),$ (1)
where the attractive component is influenced by the current destination for the target drone $n$ (this destination is changed once the drone reaches it); the repulsive one is effected by obstacle drone $j$, $j\in\mathcal{N}^{-}$, where $\mathcal{N}^{-}$ defines the set of drones except the target drone $n$. Then, the attractive component can be formulated as:
@@
However, there is a the common problem with repulsive PFG: agents stand-off in situations where they continuously repel each other without making progress, leading to a deadlock scenario [5]. To mitigate the lock, the priorities of drones are assigned randomly. Drones with higher priority exert stronger and more extensive repulsive forces, pushing lower-priority drones out of their paths. As a result, drones can reach their respective destinations in sequence without getting obstructed by other drones. The maximum radius of repulsion effect of a drone is defined as follows:
@@
where $P_{j}$ denotes the priority index of obstacle drone $j$. The higher value of $P_{j}$ is, the higher priority of a drone is. $D^{\mathsf{min}}$ denotes the minimum distance between drones before they collide. This paper sets 25cm considering the wind force caused by the Crazyfly. Thus, we can formulate the update of a vector in repulsive PFG $V^{\mathsf{r}}_{i,j}$ with index of $i$ per timestamp $t$ as follows:
@@
where $D_{i,j}$ indicates the distance between the vector $i$ and an obstacle drone $j$; $S_{j}$ denotes the scaling factor for the strength of repulsion, higher than the strength (or magnitude) of attractive component of the vector $V^{\mathsf{a}}_{i}$. This ensures that the repulsive force acting on the target drone is stronger than attractive force, thus influencing the vector $V_{i}$ according to Eq.(1). We set the scaling value $\delta=2.5$ empirically to ensure that the lowest priority drone is strong enough to repel the other drones under the attractive forces. Therefore, the vectors in PFG prioritize maintaining a safe distance and prevent potential collisions over reaching the destinations. For normalization, we set the magnitude of both repulsive and attractive components as $|V^{\mathsf{r}}_{i}(1)|=|V^{\mathsf{a}}_{i}(1)|=1$. Finally, the target drone is forced by the summation of all vectors per timestamp, formulated as $\sum^{I}_{i=1}V_{i}(t)$ for $I$ vectors.

### v2

@@ III-A4 Collision detection and collision avoidance
It detects all possible in-flight and static collisions, i.e., intersections of flight paths or walls in multi-drone missions, and then minimizes the likelihood of collisions during the path planning. This requires an intelligent path planning algorithm, making distributed sensing cost-effective and safe within various scenarios. The algorithm detects intersections of flight paths and obstacles, and augments the navigation and sensing plans selected via swarm intelligence to prevent the potential collisions, minimizing in-flight risks, i.e., the traveling distance at high risk of collisions. Fig 1 illustrates three typical types of collisions considered in this paper (cross, parallel, and destination-occupied).
@@ III-B Architecture overview
Fig. 2 illustrates an overview of M-SET architecture. The core of M-SET lies in two software approaches. One is the decentralized multi-agent collective learning method, named Economic Planning and Optimized Selections (EPOS) [6, 7, 22]. It generates for each agent a finite number of discrete navigation and sensing options, each with an estimated power consumption: the possible plans and their cost respectively. Plan generation is performed using the drones specifications (weight, propeller and battery parameters), and the sensing environment (wind and grid cells). Then the agents interact iteratively in a bottom-up and top-down fashion over a tree communication structure to make a selection such that all choices together add up to maximize the sensing quality (matching). However, these selected plans have no information about the potential collisions, and thus the path planning algorithm for collision avoidance is required to augment selected plans. This testbed uses the artificial potential field algorithm [5] to generate attractive forces towards the destinations and repulsive forces between drones. These forces guide the drones to their target cells for sensing according to the selected plan, while also avoiding potential collisions. Finally, the algorithm produces the output plans containing the X/Y coordinates and sensing (hovering) duration at each cell. They are then executed by drones in hardware, which records the energy consumption, the collected sensor data and collision risk distance during the sensing mission. Therein, collision risk distance implies the specific distance at which the risk of collisions becomes significant, illustrated by the path length within the shaded area shown in Fig. 1.
@@ IV-A Hardware UAVs for sensing
M-SET uses the Crazyflie 2.1 because of its size, weight and accessibility [^3] https://www.bitcraze.io/products/crazyflie-2-1/. The drone has 27g weight, 47mm propeller length, a battery capacity of 250mAh (LiPo battery) and it is configured to fly with an average ground speed of 0.1m/s. Based on this information, the power consumption can be estimated. Crazyflie can be programmed in Python with the support of API in Bitcraze [^4] https://github.com/bitcraze/crazyflie-lib-python. This is further expanded by the Crazyswarm API [^5] https://github.com/USC-ACTLab/crazyswarm, which provides additional functionality to control a swarm of Crazyflies. The drone can be mounted by multiple hardware decks to support different functions, such as positioning (with the support of lighthouse base stations), camera recording, and wireless charging. Due to maximum weight limitation of Crazyflie, two types of drones are assembled, one is for testing wireless charging, see Fig. 3(a), and the other is for collecting sensor data, see Fig. 3(b). The data-collection drone with an ultra low power 320×320 grayscale camera is equipped with a tiny mirror to take videos of ground.
@@ IV-B Setting up an indoor sensing environment
A 75 inch screen is set on the ground as an indoor environment to emulate the outdoor sensing environments. As shown in Fig. 4(a) and Fig. 4(b), the screen is divided into 2x3 square grid cells, each with an area of 55x47cm. Each cell can be sensed effectively from an altitude of 50cm (i.e., the vehicles can be clearly observed). Other accessory equipment is set, including lighthouse base stations and wireless chargers. To show the significant and broad impact of M-SET on a transportation scenario, the screen displays the video of the traffic flow of vehicles recorded from satellites or other simulators, e.g., Simulations of Urban Mobility (SUMO). In this scenario, a higher sensing requirement in a cell represents a high urgency for traffic monitoring over this area, which requires drones to hover a longer time to measure accurately. M-SET uses a real-world dataset of vehicle trajectories named pNEUMA [^6] https://open-traffic.epfl.ch/ collected by a swarm of drones in the congested downtown area of Athens, Greece [8].
@@ IV-C Swarm intelligence using collective learning
Figure 5: An example of collision avoidance using artificial potential field: (a) The blue drone is the target drone attracted by the destination, whereas the red drone is the obstacle drone repelling the target drone. The vectors influenced by both attractive and repulsive forces point towards the navigation of the target drone. (b) Since drone 2 has higher priority and stronger repulsive force than drone 1, drone 1 is pushed out of the target cell and "wait" until drone 2 passes by, and makes a "turn" when traveling to the next cell.
@@ IV-D Collision avoidance using artificial potential field
M-SET applies an artificial potential field algorithm [5] to path planning of drones, which creates force to repel drones from obstacles and attract them towards their destinations. As shown in Fig. 5, a Potential Fields Grid (PFG) in this implementation is created for each drone (named as the target drone) per timestamp. It is a 2D-grid of vectors, where each vector points in the direction that the target drone should fly at that position per timestamp [5]. There are two types of PFG: attractive PFG and repulsive PFG, which coexist to balance the drone navigation. The attractive PFG generates forces that pull the drone towards its target destination. The repulsive PFG creates forces that push the drone away from obstacles (e.g., the other drones, walls and other static obstacles). These repulsive forces are stronger when the target drone is closer to an obstacle. After the summation of all types of forces, the drone navigates towards its destination while avoiding obstacles. Thus, given $I$ vectors, a vector with index of $i$ at timestamp $t$, $i\geq 1$, $t\geq 1$, consists of two components: attractive component $V^{\mathsf{a}}$ and repulsive component $V^{\mathsf{r}}$, formulated as follows:
$V_{i}(t)=\sum_{j\in\mathcal{O}}V^{\mathsf{r}}_{i,j}(t)+V^{\mathsf{a}}_{i}(t),$ (1)
where the repulsive one is effected by the obstacle $j$, $j\in\mathcal{O}$, where $\mathcal{O}$ defines the set of obstacles in the map; the attractive component is influenced by the current destination for the target drone $n$ (this destination is changed once the drone reaches it). The attractive component can be formulated as:
@@
However, there is a the common problem with repulsive PFG: agents stand-off in situations where they continuously repel each other without making progress, leading to a deadlock scenario [5]. To mitigate the lock, the priorities of drones are assigned randomly. Drones with higher priority exert stronger and more extensive repulsive forces, pushing lower-priority drones out of their paths. As a result, drones can reach their respective destinations in sequence without getting obstructed by obstacles. The maximum radius of repulsion effect of a drone is defined as follows:
@@
where $P_{j}$ denotes the priority index of the obstacle $j$. The higher value of $P_{j}$ is, the higher priority of an obstacle drone is ($P_{j}$ is a constant if $j$ is a static obstacle like walls). $D^{\mathsf{min}}$ denotes the minimum distance between a drone and an obstacle before they collide. This paper sets 25cm considering the wind force caused by the Crazyfly. Thus, we can formulate the update of a vector in repulsive PFG $V^{\mathsf{r}}_{i,j}$ with index of $i$ per timestamp $t$ as follows:
@@
where $D_{i,j}$ indicates the distance between the vector $i$ and an obstacle $j$; $S_{j}$ denotes the scaling factor for the strength of repulsion, higher than the strength (or magnitude) of attractive component of the vector $V^{\mathsf{a}}_{i}$. This ensures that the repulsive force acting on the target drone is stronger than attractive force, thus influencing the vector $V_{i}$ according to Eq.(1). We set the scaling value $\delta=2.5$ empirically to ensure that the lowest priority drone is strong enough to repel the other drones under the attractive forces. Therefore, the vectors in PFG prioritize maintaining a safe distance and prevent potential collisions over reaching the destinations. For normalization, we set the magnitude of both repulsive and attractive components as $|V^{\mathsf{r}}_{i}(1)|=|V^{\mathsf{a}}_{i}(1)|=1$. Finally, the target drone is forced by the summation of all vectors per timestamp, formulated as $\sum^{I}_{i=1}V_{i}(t)$ for $I$ vectors.